OpenStudy (anonymous):

compute: 7 + [- 3/7 ]

OpenStudy (ajprincess):

\((+)*(-)=(-)\) so \[7+(-\frac{3}{7})\] \[=7-\frac{3}{7}\] \[=\frac{7}{1}-\frac{3}{7}\] The l.c.m is 7 so \[=\frac{7}{1}-\frac{3}{7}\] \[=\frac{7*7}{1*7}-\frac{3}{7}\] \[=\frac{(7*7)-3}{7}\] Hope u can do nw:)

OpenStudy (anonymous):

46/7

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

ok i was writing it wrong once i multiplied 7*7

OpenStudy (anonymous):

your break down helped

OpenStudy (ajprincess):

:)